Given the linear correlation coefficient r and the sample size n, determine the critical values of r and use your finding to state whether or not the given r represents a significant linear correlation. Use a significance level of 0.05. r = 0.353, n = 15
Answer #1

To answer this question, I used the Critical Values of the Pearson Product-Moment Correlation Coefficient table.

The level of significance for a two-tailed test is 0.05

The degrees of freedom = n - 2 = 15 - 2 = 13

(from the table) The critical value of the correlation coefficient is ±0.514

The value of r given in the problem statement (r = 0.353) is less than 0..514 This shows there is no significant linear correlation.